OpenStudy (anonymous):

-6, 7, 8 -16, 32, 36 -5, 12, 13 -8, 12, 16 State which set of three numbers could represent the lengths of the sides of a right triangle. please help me

OpenStudy (cwrw238):

lengths cannot be negative

OpenStudy (texaschic101):

pythagorean theorem

OpenStudy (anonymous):

They're not negative, I just used dashes as an alternative to letters. My bad lol.

OpenStudy (cwrw238):

ok

OpenStudy (anonymous):

Yes, so just go through each of them, and do exactly what you did in the last problem.

OpenStudy (cwrw238):

as texaschic said see which fit the pythagoras theorem

OpenStudy (anonymous):

Use Pythagorean theorem on each of them until you find one where $$ A^2 + B^2 = C^2$$

OpenStudy (texaschic101):

just make sure that the biggest number is C, the hypotenuse

OpenStudy (anonymous):

Not sure what it's supposed to equal or what I'm supposed to be looking for.

OpenStudy (texaschic101):

example : A^2 + B^2 = C^2 ---- (6,7,8) 6^2 + 7^2 = 8^2 36 + 49 = 64 85 = 64....Nope...this one is not

OpenStudy (anonymous):

O wait

OpenStudy (anonymous):

I think I get it

OpenStudy (texaschic101):

just make sure that the biggest number is in the C position

OpenStudy (anonymous):

Okay thank youuuu

OpenStudy (texaschic101):

no problem :)